# Approaching Rate-Distortion Limits in Neural Compression with Lattice Transform Coding

### v1

@@ Appendix A Experimental Details
The code for these experiments can be found at https://github.com/leieric/lattice-transform-coding. In all experiments, we use a batch size of 64 and train until convergence using Adam. For the i.i.d. scalar sequences, we found that using CELU nonlinearities with no biases in the transforms sometimes helped improve training stability.
Source generation.
The Gaussian and Laplacian sources are straightforward to simulate. The Banana source is taken from the Tensorflow Compression library [Ballé et al., 2024]. To generate the Banana-1d marginal, we simply take the first dimension of the 2-d Banana realizations. The Sawbridge source [Wagner and Ballé, 2021] is a continuous random process
$X(t)=t-\mathbbm{1}\{t\geq U\},\quad 0\leq t\leq 1,$ (16)
where $U\sim\mathrm{Unif}([0,1])$. Following Wagner and Ballé [2021], we uniformly sample 1024 time steps between 0 and 1, which creates 1024-dimensional vectors. For these sources, we sample $10^{7}$ samples as our training dataset. Finally, the Physics and Speech sources are taken from https://github.com/mandt-lab/RD-sandwich.
Efficiency.
The main computational bottleneck that LTC has is the use of Monte-Carlo integration for (11). Sampling uniformly random vectors from the Voronoi region of a lattice requires a quantization step, and hence solving the CVP. Since we use 4096 sampled vectors per batch, this means CVP needs to be solved on all 4096 vectors. The lattices chosen above all have GPU-friendly implementations that can parallelize across the input vectors. On NVIDIA RTX 5000 GPUs with 16 GB memory, LTC training until convergence took at most a few hours for the Speech and Physics sources, and minutes for the i.i.d. scalar sequences.
Appendix B Additional Figures
Monte-Carlo Samples.
For lattices of lower dimensions, we found that setting the number of Monte-Carlo samples $N_{\text{int}}$ for approximation of (11) to be 4096 was sufficient to recover near-ECVQ performance. For the Leech lattice of dimension 24, however, better performance could be recovered by increasing $N_{\text{int}}$ at the cost of increased GPU memory requirements; see Fig. 12.
@@ Appendix B Additional Figures
Figure 13: NTC quantization regions in the latent space (integer lattice).
Figure 14: Rate-redundancy of Laplacian i.i.d. sequences with L1 distortion.
Figure 15: Rate-redundancy of Sawbridge i.i.d. sequences with L2 distortion. Here, we measure rate-redundancy with respect to $H(D)$, since a closed-form expression for $H(D)$ is known, but not for $R(D)$.
Figure 16: Ablation study on entropy models. Cross-entropy (CE) is plotted for both factorized prior (FP) and normalizing flow (Flow) density models, as well as the true entropy of the quantized latent $H(\hat{\bm{y}})$. The cross-entropy from the normalizing flow is a much tighter upper bound on $H(\hat{\bm{y}})$ compared to that of the factorized density.
Figure 17: One-shot compression of the Physics source.
Figure 18: Performance of BLTC on i.i.d. sequences of Physics realizations, where $n$ is the block-length (number of realizations compressed).

### v2

@@ Appendix A Experimental Details
The code for these experiments will be released upon publication. In all experiments, we use a batch size of 64 and train until convergence using Adam. For the i.i.d. scalar sequences, we found that using CELU nonlinearities with no biases in the transforms sometimes helped improve training stability. For synthetic sources, Speech and Physics, the rate-distortion Lagrange multiplier $\lambda$ is swept over 0.5, 1, 1.5, 2, 4, 8. For images, we use the default $\lambda$ values in Bégaint et al. (2020). In addition, for images, a product lattice is applied along the channel dimension of the latent tensor, which has shape $C\times H\times W$. There are $C/n$ lattices applied $H\cdot W$ times, where $n$ is the lattice dimension.
A.1 CVP Algorithms
In this paper, we use the CVP algorithms outlined in Conway & Sloane (1999, Ch. 20). For the Barnes-Wall lattice $\Lambda_{16}$ in dimension 16, we use the algorithm outlined in Conway & Sloane (1984). Finally, the Leech lattice $\Lambda_{24}$ in dimension 24 has the slower method described in Conway & Sloane (1984) and the improved version in Conway & Sloane (1986). In many cases, the lattice $\Lambda=\bigcup_{i=1}^{T}(\bm{r}_{i}+\Lambda_{0})$ is a union of $T$ cosets of a sublattice $\Lambda_{0}$, and one can solve CVP for $\Lambda$ by computing $Q_{\Lambda}(\bm{x})=\operatorname*{arg\,min}_{1\leq i\leq T}Q_{\Lambda_{0}}(\bm{x}-\bm{r}_{i})+\bm{r}_{i}$. For example, the $E_{8}$ lattice is equivalent to two cosets of $D_{8}$.
A.2 Monte-Carlo Samples
Regarding variable-rate LTC, for lattices of lower dimensions, we found that setting the number of Monte-Carlo samples $N_{\text{int}}$ for approximation of (5) to be 4096 was sufficient to recover near-ECVQ performance. For the Leech lattice of dimension 24, however, better performance could be recovered by increasing $N_{\text{int}}$ at the cost of increased GPU memory requirements; see Fig. 10.
@@ A.2 Monte-Carlo Samples
A.3 Source generation
The Gaussian and Laplacian sources are straightforward to simulate. The Banana source is taken from the Tensorflow Compression library (Ballé et al., 2024). To generate the Banana-1d marginal, we simply take the first dimension of the 2-d Banana realizations. The Sawbridge source (Wagner & Ballé, 2021) is a continuous random process
$X(t)=t-\mathbbm{1}\{t\geq U\},\quad 0\leq t\leq 1,$ (10)
where $U\sim\mathrm{Unif}([0,1])$. Following Wagner & Ballé (2021), we uniformly sample 1024 time steps between 0 and 1, which creates 1024-dimensional vectors. For all synthetic sources, we sample $10^{7}$ samples as our training dataset. Finally, the Physics and Speech sources are taken from https://github.com/mandt-lab/RD-sandwich.
A.4 Efficiency
Compared to NTC, LTC requires more memory/training time, roughly 50% more for the low/moderate dimension sources and 10% more for images. The main computational bottleneck that LTC has is the use of Monte-Carlo integration for (5). Sampling uniformly random vectors from the Voronoi region of a lattice requires a quantization step, and hence solving the CVP. Since we use 4096 sampled vectors per batch, this means CVP needs to be solved on all 4096 vectors. The lattices chosen above all have GPU-friendly implementations that can parallelize across the input vectors. On NVIDIA RTX 5000 GPUs with 16 GB memory, LTC training until convergence took at most a few hours for the Speech and Physics sources, and minutes for the i.i.d. scalar sequences. This was roughly twice the amount of time to train NTC. However, on large-scale images, the training speed is dominated by the backwards pass through the convolutional neural networks, and so LTC takes around the same time as NTC (around 10 days per rate point).
Appendix B Additional Results
BD-rate and runtime tables for Kodak.
Here, we show the BD-rate tables (Tab. 1) and full runtime comparison (Tab. 2) on the Kodak benchmark. For the runtime comparison, it is time taken per-image on a single-core CPU, averaged over the rate-distortion tradeoff. We see that Cheng2020-LTC-$\Lambda_{24}$ outperforms all other methods in rate-distortion performance, including ELIC (He et al., 2022). We note that the ELIC model is NTC-based, and thus even further gains can be expected by integrating LTC with ELIC transforms and entropy model. This would also inherit the efficient nature of the ELIC layers, as the quantization and entropy modelling are independent components of the neural compressor. We leave this for future work as it is out of the scope of the current study. Regarding runtime on the Cheng2020 architecture, using the $E_{8}$ lattice incurs a 1.23x increase in runtime over NTC, whereas the Leech lattice incurs a 2.68x increase. We note that the Leech lattice CVP solver we used (Conway & Sloane, 1986) is not the fastest; one can expect faster runtimes with the Leech decoder of Vardy & Be’ery (1993).
Table 1: BD-rate comparison on Kodak.
Model BD-rate (NTC anchor) BD-rate (VTM anchor)
Cheng2020-NTC 0.000 5.210
Cheng2020-LTC-$D_{n}^{*}$2.337 2.753
Cheng2020-LTC-$E_{8}$5.274 -0.335
Cheng2020-LTC-$\Lambda_{24}$16.708 -11.895
NVTC (Feng et al., 2023) -4.953 -2.919
VTM (Bross et al., 2021) -7.879 0.000
ELIC (He et al., 2022) -10.607 -5.656
Table 2: Inference encoding runtime on Kodak, single CPU. LTC models use 1024 samples for Monte-Carlo estimation.
Model Runtime (sec.)
Cheng2020-NTC 3.241
Cheng2020-LTC-$D_{n}^{*}$ 4.016
Cheng2020-LTC-$E_{8}$ 4.002
Cheng2020-LTC-$\Lambda_{24}$ 8.634
ELIC (He et al., 2022) 3.132
NTC with lattice quantization during inference.
Here, we show the performance of LTC when the transforms and entropy model are loaded from a pre-trained NTC model. In other words, a trained NTC model’s scalar quantizer is replaced by a lattice quantizer at inference time. The performance of this scheme is largely dependent on the source at hand. Fig. 12 shows that on the Gaussian source, this scheme performs the same as training LTC from scratch; it may shift the point on the R-D tradeoff but overall lies on the same R-D curve. This is perhaps not unexpected, as the transforms learned by NTC and LTC for the i.i.d. Gaussian source effectively perform a scaling of the source, and so interchanging them has no effect. We use the same Flow entropy model for fair comparison. In general, however, this is not expected to be the case, as the analysis transform determines which region of the lattice gets used; for more complex sources, there may be more sophisticated ways the transform determines how the lattice is used. For example, in Fig. 12, we demonstrate the same scheme on Kodak images with the Cheng2020-LTC-$\Lambda_{24}$ model. Here, we see that while using a Leech lattice quantizer with the trained NTC model improves the performance over NTC by a BD-rate of -12.984%, training the model with the lattice quantizer helps boost the performance further to a BD-rate of -16.708%. This result implies that one can avoid training from scratch when a good set of transforms and entropy model are already trained from a NTC architecture, and instead potentially fine-tune the NTC transforms and entropy model with the chosen lattice for best performance. In the future, when better transforms and entropy models are developed, one can also train with the lattice from scratch to get the best performance, as the training times between NTC and LTC (from scratch) are on the same scale (weeks).
Figure 11: Gaussian source, comparing LTC (trained) versus LTC loaded with pretrained NTC weights.
Figure 12: Kodak images, comparing LTC (trained) versus LTC loaded with pretrained NTC weights.
Learned quantizer cells in regions of low probability.
In Figs. 13, 14, we visualize the quantizer cells of LTC on regions of low probability density, for the 2D Gaussian source and Banana source, respectively. That is, we train LTC on a source $p_{\bm{x}}$, and evaluate on the support of $p_{\bm{x}}$ as well as regions outside the support of $p_{\bm{x}}$. Overall, the quantization cells successfully generalize outside the support of $p_{\bm{x}}$, meaning that if a sample $\bm{x}$ to be compressed is perturbed slightly outside the support of $p_{\bm{x}}$, its reconstruction will still be close to $\bm{x}$.
(a) In-distribution samples.
(b) Out-of-distribution samples.
Figure 13: Quantizer cells of LTC trained on 2-d i.i.d. Gaussians.
(a) In-distribution samples.
(b) Out-of-distribution samples.
Figure 14: Quantizer cells of LTC trained on Banana source.
Rate-redundancy.
For certain figures in the ablation study, we define the normalized rate-redundancy of a compressor $\mathsf{c}$,
$\mathcal{E}_{\mathsf{c}}(D):=({R_{\mathsf{c}}(D)-R(D)})/{R(D)},$ (11)
where $R_{\mathsf{c}}(D)$ is the rate achieved by $\mathsf{c}$ at distortion $D$, and $R(D)$ is the rate-distortion function of $S$. Intuitively, $\mathcal{E}_{\mathsf{c}}(D)$ measures the fraction of additional bits per sample required to code a sequence relative to $R(D)$, and makes some of the performance gains easier to see in the figures.
(a) Laplace source ($L_{1}$).
(b) Banana-1D ($L_{2}$).
Figure 15: Rate-distortion performance of scalar i.i.d. sequences.
(c) LTC, sub-opt lattices.
(d) Dither vs. STE.
Figure 16: Ablation study, i.i.d. Gaussians.
(a) NTC (integer lattice).
(b) LTC (hexagonal lattice).
Figure 17: Latent space quantization regions.
Figure 18: Flow provides tighter entropy estimate.
(a) $R\approx 3$.
(b) $R\approx 5$.
Figure 19: ECVQ quantization regions, banana source.
(c) $R\approx 3$.
(d) $R\approx 5$.
Figure 20: NTC quantization regions, Banana source.
(a) $R\approx 3$.
(b) $R\approx 5$.
Figure 21: LTC quantization regions, Banana source.
(a) $R\approx 3$.
(b) $R\approx 5$.
Figure 22: Quantization regions of Banana source using ECLQ. Without the transforms, quantization regions of ECLQ fail to conform to the source manifold.
B.1 Ablation Study
We first provide a comparison with entropy-constrained lattice quantization (ECLQ), which applies lattice quantization directly to the source without transforms. While ECLQ approaches VQ performance in the large rate regime, it is highly sub-optimal in the low/moderate rate regime. LTC, on the other hand, performs much better at all rates. This is shown in Fig. 23 for Gaussians and Banana. While the quantization regions look similar, the primary benefit of LTC is that the $g_{s}$ transform moves the reconstructions $\hat{\bm{x}}$ according to the source statistics. Standard ECLQ keeps the reconstructions at the center of the lattice cells, which is sub-optimal for non-uniform sources. Moreover, on the Banana source, without the transforms, ECLQ fails to match the quantizer cells to the source statistics (Fig. 22), unlike LTC (Fig. 21). We further analyze (i) the training objective used, and (ii) the effect of density models on the entropy estimate of $\hat{\bm{y}}$. For the training objective used, we compare using dithered quantization with STE for the rate term (i.e., using likelihoods from (6) and (5) respectively). Shown in Fig. 15(d), for low rates, the reported cross-entropies from dithering are much larger than that of STE, but converge to the STE performance as rate increases. One possible reason is that the convolved density in (6) over-smooths the true density of $\bm{y}$ in the low-rate regime, leading to sub-optimality.
(a) i.i.d. Gaussian.
(b) Banana source.
Figure 23: LTC vs. ECLQ.